# Scalp Diagnostic System With Label-Free Segmentation and Training-Free Image Translation

### v2

@@ Appendix A Detailed Overview of Scalp Diseases
Appendix B Implementation Details
Hair segmentation. For our heuristic-driven hair segmentation, we utilized the U2-Net [39] official source code to train $3{,}000$ pseudo image-label pairs. The training parameters included a batch size of $32$, a constant learning rate of $0.001$, $100$ training epochs, and the Adam optimizer. The output from U2-Net was binarized to obtain $\hat{M}$, using a threshold of $0.5$.
@@ Appendix B Implementation Details
Image augmentation. For DiffuseIT-M, detailed in Section 3.3 of the main paper, we construct a loss function for image translation while preserving hair content. This function is defined as follows:
@@
$\begin{gathered}\ell_{sty}(x_{trg},\hat{x}_{0}(x_{t}))=||\mathbf{c}(x_{trg})-\mathbf{c}(\bar{x})||_{2}\\ +\lambda_{mse}||x_{trg}-\hat{x}_{0}(x_{t})||_{2},\end{gathered}$ (19)
where $\lambda_{mse}$ is set to $3{,}000$, and the weight for $\ell_{style}$, $\lambda_{1}$, is set to $2{,}000$.
@@
with $cos_{ij}(x)$ representing the cosine distance between $k^{l}_{i}(x)$ and $k^{l}_{j}(x)$. The weights $\lambda_{sim}$ and $\lambda_{con}$ are set to $1{,}000$ and $200$, respectively. Additionally, weights $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to $1,000$, $100$, and $200$, respectively. During our experiment, the model is configured to generate images with a resolution of $256\times 256$ pixels, utilizing a denoising step of $1{,}000$.
For implementation previous studies, we have performed the official source code of both DiffuseIT [27] and AGG [28], conducting experiments with the provided hyperparameters.
@@
Alopecia severity prediction. To predict the severity of alopecia, we focused on extracting hair-related features, specifically hair thickness and count. For this, we employed the OpenCV library, utilizing Non-Maximum Suppression (NMS), Morphology operation, and Hough transformation techniques. These features were then used to train a prediction model with the GradientBoosting Classifier, a component of the Scikit-learn library [5]. The model’s training parameters were set as follows: a cross-entropy loss function, a learning rate of $0.01$, a maximum depth of $5$, $3{,}000$ estimators, and a subsample ratio of $1$. All other parameters remained at their default settings in Scikit-learn.
In comparison, for the image-based models, we trained DenseNet169 and EfficientFormerV2 using similar hyperparameters as those used for scalp disease classification. The key differences were the employment of the cross-entropy loss function and the use of the Adam optimizer for these models.
Appendix C Pseudo Image and Mask Visualization
@@
Appendix D Scalp Disease and Severity Classification
@@
Input: $severities$: A collection of records for each severity. Assume that each element of $severities$ is classified as 0 (good), 1 (mild), 2 (moderate), and 3 (severe).
Output: $ratios$: Sampling ratio among four severity levels.
1:  $\epsilon\leftarrow 1\times 10^{-9}$
[…]
3:  for each $sevLevel$ in $[0,1,2,3]$ do
4:     $sevCounts[sevLevel]$leftarrow 00
5:  end for
6:  for each $severity$ in $severities$ do
7:     $sevCounts[severity]$leftarrow $sevCounts[severity]+1$
8:  end for
[…]
10:  for each $sevLevel$ in $[0,1,2,3]$ do
11:     $invCounts[sevLevel]$leftarrow $1/(sevCounts[sevLevel]+\epsilon)$
12:  end for
13:  $normFactor$leftarrow $\sum_{sevLevel=0}^{3}invCounts[sevLevel]$
14:  $ratios$leftarrow [][][ ]
15:  for each $sevLevel$ in $[0,1,2,3]$ do
16:     $ratio\leftarrow invCounts[sevLevel]/normFactor$
17:     $ratios[sevLevel]$leftarrow $ratio$
18:  end for
19:  return $ratios$
D.1 Data Augmentation Strategy
Figure 12: Data distribution of AI-Hub scalp dataset before and after the application of augmentation techniques.
@@ D.1 Data Augmentation Strategy
Figure 12 displays the data distribution before and after this augmentation. The post-augmentation distribution shows a more balanced representation across various classes, especially a rise in the severe category and a reduction in the good category. This balanced dataset played a crucial role in enhancing our model’s performance by providing an even distribution of the training samples. The same augmentation strategy was applied across all diffusion-based methods.
D.2 Impact of Backbone Model
We assessed the performance of scalp disease classification using different pretrained backbone models. Our evaluation metric was the F1 macro score, which we also used to compare our results with existing augmentation methods, including those utilizing DiffuseIT and AGG. The experiments involved two CNN-based models (ResNet [16] and ResNeXt [51]) and two Transformer-based models (ViT [12] and EfficientFormer [29]), all maintaining consistent hyperparameters as described in Section 11.
@@ D.2 Impact of Backbone Model
Figure 13: Qualitative results of hair count estimation. The displayed lines on the masks represent the line segments identified through the Hough Transformation.
Figure 14: Qualitative results and estimated features of hair thickness estimation. The red dots on the image indicate the intersection points of hair strands and perpendicular line of hair direction, while the green lines represents the perpendicular lines. Each caption represents severity of alopecia.
Appendix E Alopecia Severity Prediction
E.1 Feature Selection for Severity Estimation
We identified key features for alopecia severity prediction, focusing on hair thickness and count. For thickness, we analyzed the length of list $B$ and computed the mean and median values of $\mathcal{T}$, excluding outliers identified using the interquartile range (IQR). We also considered the mean and median of the entire $\mathcal{T}$. Additionally, the number of line segments detected by the hair count estimator was used as a supplementary feature.
E.2 Visualization of Severity Features
Our method estimates hair thickness using mask images. Figure 13 showcases the qualitative results of our multi-scale Hough transform, which effectively distinguishes individual hairlines in the mask images. In Figure 14, we demonstrate thickness estimation across various scenarios for four alopecia severity classes: good, mild, moderate, and severe. Our segmentation successfully identified different hair thicknesses and distinguished hair from severe scalp conditions like dandruff or erythema. Notably, as alopecia severity increased, the feature values related to hair thickness decreased, validating the effectiveness of our approach.
Appendix F Qualitative Results of DIffuseIT-M

### v3

@@ Appendix 0.A Detailed Overview of Scalp Diseases
Appendix 0.B Implementation Details
Hair segmentation. For our heuristic-driven hair segmentation, we utilized the U2-Net [25] official source code to train 3,0003{,}000 pseudo image-label pairs. The training parameters included a batch size of $32$, a constant learning rate of $0.001$, $100$ training epochs, and the Adam optimizer. The output from U2-Net was binarized to obtain $\hat{M}$, using a threshold of $0.5$.
@@ Appendix 0.B Implementation Details
Image augmentation. For DiffuseIT-M, we construct a loss function for image translation while preserving hair content. This function is defined as follows:
@@
[…]
where $\lambda_{mse}$ is set to 3,0003{,}000, and the weight for $\ell_{style}$, $\lambda_{1}$, is set to 2,0002{,}000.
@@
with $cos_{ij}(x)$ representing the cosine distance between $k^{l}_{i}(x)$ and $k^{l}_{j}(x)$. The weights $\lambda_{sim}$ and $\lambda_{con}$ are set to 1,0001{,}000 and $200$, respectively. Additionally, weights $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to $1,000$, $100$, and $200$, respectively. During our experiment, the model is configured to generate images with a resolution of $256\times 256$ pixels, utilizing a denoising step of 1,0001{,}000.
@@
Appendix 0.C Pseudo Image and Mask Visualization
@@
Appendix 0.D Scalp Disease and Severity Classification
@@
1 Input: severitiesseverities: A collection of records for each severity. Assume that each element of severitiesseverities is classified as 0 (good), 1 (mild), 2 (moderate), and 3 (severe).
2Output: ratiosratios: Sampling ratio among four severity levels.
31:  $\epsilon\leftarrow 1\times 10^{-9}$
2:  sevCountssevCounts $\leftarrow$ $\{\}$
3:  for each sevLevelsevLevel in $[0,1,2,3]$ do
4:   $sevCounts[sevLevel]$ $\leftarrow$ 0
45:  end for
6:  for each severityseverity in severitiesseverities do
57:   $sevCounts[severity]$ $\leftarrow$ $sevCounts[severity]+1$
8:  end for
69:  invCountsinvCounts $\leftarrow$ $\{\}$
10:  for each sevLevelsevLevel in $[0,1,2,3]$ do
11:   $invCounts[sevLevel]$ $\leftarrow$ $1/(sevCounts[sevLevel]+\epsilon)$
712:  end for
813:  normFactornormFactor $\leftarrow$ $\sum_{sevLevel=0}^{3}invCounts[sevLevel]$
14:  ratiosratios $\leftarrow$ $[]$
15:  for each sevLevelsevLevel in $[0,1,2,3]$ do
16:   $ratio\leftarrow invCounts[sevLevel]/normFactor$
17:   $ratios[sevLevel]$ $\leftarrow$ ratioratio
18:  end for
19:  return ratiosratios
@@
Figure 7: Distribution of data augmented using DiffuseIT-M. For comparison, the original data distribution is presented in Figure 2.
0.D.1 Data Augmentation Strategy
@@
Figure 7 displays the data distribution after this augmentation. The post-augmentation distribution shows a more balanced representation across various classes, especially a rise in the severe category and a reduction in the good category. This balanced dataset played a crucial role in enhancing our model’s performance by providing an even distribution of the training samples. The same augmentation strategy was applied across all diffusion-based methods.
@@
Impact of backbone model. We assessed the performance of scalp disease classification using different pretrained backbone models. Our evaluation metric was the F1 macro score, which we also used to compare our results with existing augmentation methods, including those utilizing DiffuseIT and AGG. The experiments involved two CNN-based models (ResNet [9] and ResNeXt [31]) and two Transformer-based models (ViT [7] and EfficientFormer [19]), all maintaining consistent hyperparameters as described in Section 6.
Appendix 0.E Qualitative Results of DIffuseIT-M